OpenStudy (anonymous):

The sum of the reciprocals of two even consecutive even numbers is 7\24. Find the numbers. I did 1 over x + 1 over x = 7\24, but that didn't work. What did I do wrong?

OpenStudy (mathstudent55):

Even numbers are x and x + 2, not x and x + 1

OpenStudy (campbell_st):

well let the consecutive numbers be x and x + 2 the reciprocals are 1/x and 1/(x+2) so you now have an equation \[\frac{1}{x} + \frac{1}{x +2} = \frac{7}{24}\]

OpenStudy (mathstudent55):

Luis, the problem states " two even consecutive even numbers " You are making the same mistake the OP made.

OpenStudy (anonymous):

do I have to make the denominators of each number the same?

OpenStudy (anonymous):

or would I multiply the first number by x and the other number by x + 2 and then multiply 7\24 by x + x + 2?

OpenStudy (mathstudent55):

"doent matter, ;se expresa de la misma forma !!" x and x + 2 is not the same as x and x + 1

OpenStudy (mathstudent55):

Get the LCD of the 3 denominators and multiply all fractions by the LCD. This will eliminate all denominators from the equation.

OpenStudy (mathstudent55):